OpenStudy (jenessa_2413):

Which of the following expressions are equivalent to −3.5(2 − 3n) − 2.5n? A. −7 − 8n B. −7 + 8n C. −7 − 13n D. −7 − n(10.5 − 2.5) E. −7 + n(10.5 − 2.5)

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

i made two typos \[ −3.5(2 − 3n) − 2.5n\] multiply out get \[-3.5\times 2-3.5\times 3-2.5n\]

OpenStudy (misty1212):

dang another typo \[-3.5\times 2-3.5\times 3n-2.5n\]

OpenStudy (misty1212):

now you got \[-7+10.5n-2.5n\] last step is to combine like terms

OpenStudy (misty1212):

do \(10.5-2.5\) then put the \(n\) next to it

OpenStudy (jenessa_2413):

Ok thanks so much @misty1212

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]